# Low-altitude Friendly-Jamming for Satellite-Maritime Communications via Generative AI-enabled Deep Reinforcement Learning

### v1

@@ 3.2 LEO Satellite Orbit Model
LEO satellites orbit at altitudes ranging from about 500 to 2,000 kilometers above the surface of Earth [43]. Since the orbital altitudes are relatively low, LEO satellites are characterized by fast orbital speeds, usually circling the Earth every 90 to 120 minutes [5]. Mathematically, the motion of the LEO satellite is usually described by the Keplerian six elements $\left\langle\beta,\omega,\Omega,e,a,\vartheta\right\rangle$ [44], as shown in the orbit model in Fig. 3, which is introduced in detail as follows.
@@
We consider that the LEO satellite orbit is circular [43]. In this case, $e$ and $\vartheta$ are set to 0, and $a$ is equal to the orbital radius $l_{S}$. Accordingly, the elements of the LEO satellite in $m$ orbit at $t$ time can be set as $\left\langle\beta_{m},\omega_{m}[t],\Omega_{m},{l}_{Sm}\right\rangle$, wherein $\omega_{m}[t]=\omega_{m}[t]+2\pi\left(({t}/{\mathcal{T}_{Sm}})\mod 1\right)$ and $\mathcal{T}_{Sm}$ denotes the orbital period of the LEO satellite. Moreover, $l_{Sm}=H_{Sm}+R_{E}$, where $H_{Sm}$ and $R_{E}$ are the orbital altitude of the LEO satellite and the radius of Earth, respectively. In this case, the orientation of the LEO satellite in the 3D Cartesian coordinates $(x_{Sm}[t],y_{Sm}[t],z_{Sm}[t])$ at time slot $t$ is expressed by
@@ 3.3 Vessel Movement Model
As shown in Fig. 3, two three-dimensional right-handed Cartesian coordinate systems of a vessel are generally used to represent movement of the vessel [45]. Specifically, one is a general coordinate system $\left\{g\right\}$ in which the origin is on the surface of the sea, and the defined $x$, $y$, and $z$ axes are pointing the north, east, and down, respectively. The other is a fixed coordinate system $\left\{f\right\}$ in which the origin is the gravity of the vessel, with $x^{f}$, $y^{f}$, and $z^{f}$ pointing to the bow, starboard, and down, respectively. The rotations around the $x^{f}$, $y^{f}$, and $z^{f}$ axes are defined as roll ($\phi$), pitch ($\theta$), and yaw ($\psi$), respectively. Moreover, the Euler angle vector is denoted as $\boldsymbol{\Theta}=\left[\phi,\theta,\psi\right]^{T}$. Mathematically, the six degree-of-freedom (DOF) vessel model $\left\langle x,y,z,\phi,\theta,\psi\right\rangle$ is used to represent movement of a vessel, which is as follows [46]:
@@ 3.4.1 S2V Link from LEO Satellite
During the S2V link, we consider that the vessels can detect the signals from the LEO satellite by using the approach in [28], thus obtaining the quantized form of the actual channel state information (CSI). In this case, since the altitude of the LEO satellite is sufficient for line-of-sight (LoS) transmission [47], we employ a typical composite channel to represent the communications from the LEO satellite to vessels. The channel of the S2V link at time slot $t$ can be expressed by [42]
@@
where $C_{S}$ and $W_{S}$ are the path loss parameter and exponent, respectively. Moreover, ${d_{S,V}[t]}$ is the distance between the LEO satellite and vessel (Alice or Eve) at time slot $t$, which is computed by the 3D positions of the LEO satellite and vessel according to Eqs. (1) and (2), respectively. In addition, $\delta_{S,V}[t]$ is the zero-mean Gaussian random variable with standard deviation $\sigma_{X_{S}}$ [48][49]. Note that $d_{S,A}[t]$ and $d_{S,E}[t]$ are the distance from the LEO satellite to Alice and Eve at time slot $t$, respectively, which are used to calculate corresponding path loss ($PL_{S,A}[t]$ and $PL_{S,E}[t]$), and get the matching channels ($h_{S,A}[t]$ and $h_{S,E}[t]$).
@@ 3.4.2 U2V Link from UAV
where ${d_{U,V}[t]}$ denotes the distance from the UAV to vessel (Alice or Eve) at time slot $t$, which is computed based on the movements of the UAV and vessel. Moreover, $C_{U}$ denotes the path loss parameter, $d_{c}$ denotes the reference distance, $W_{U}$ and $\delta_{U,V}[t]$ are the path loss parameter and zero-mean Gaussian random variable with $\sigma_{X_{U}}$, respectively [50]. Note that $PL_{U,A}[t]$ and $PL_{U,E}[t]$ denote the path loss from the UAV to Alice and Eve at time slot $t$, respectively, which are obtained by the distance from the UAV to Alice and Eve ($d_{U,A}[t]$ and $d_{U,E}[t]$), respectively.
@@
where $P_{{S}}$ is the transmit power of the LEO satellite, $G_{S}$ and $G_{U}$ denote the antenna gains of satellite and UAV, respectively. Moreover, $G_{S,S}$ is the antenna gain of the vessel served by the LEO satellite, $P_{{U}}[t]$ denotes the transmit power of the UAV at time slot $t$, and $\sigma^{2}$ denotes the maritime additive white Gaussian noise power.
@@
where $\left[\chi\right]^{+}$ indicates the value at which the larger of 00 and $\chi$.
@@ 3.5 UAV Energy Consumption Model
Furthermore, we introduce the moving UAV energy consumption. Commonly, the total energy consumption of a UAV is composed of communication energy and propulsion energy. However, the communication energy is usually ignored during the calculation because its value is extremely small compared to propulsion energy [51]. Thus, the propulsion power consumption of the UAV in 2D horizontal state is calculated by [51]
@@
where $v_{h}[t]=\sqrt{(a_{x}^{U}[t])^{2}+(a_{y}^{U}[t])^{2}}/t$ is the horizontal direction velocity of the UAV, $P_{I}$ and $P_{B}$ are the induced power and blade profile power, respectively, $v_{tip}$ and $v_{i}$ are the rotor blade tip velocity and average rotor induced velocity, respectively. Moreover, $r_{a}$, $s_{r}$, $a_{r}$, and $\rho$ denote the airframe drag ratio, rotor robustness, rotor disk area, and air density, respectively.
Note that we exclude the energy consumed by the UAV accelerating or decelerating, since this process occupies only a minor fraction of the total UAV flight time [52]. Therefore, we use a heuristic closed approximation to express the 3D energy consumption of the UAV, where the considerations include the propulsive energy, kinetic energy, and gravitational energy during the ascent and descent of the UAV over time. The 3D trajectory energy consumption of the UAV is denoted by [53]
@@
where $v_{f}[t]=\sqrt{\left(v_{h}[t]\right)^{2}+\left(v_{v}[t]\right)^{2}}$ is the UAV forward direction velocity at time slot $t$, of which $v_{v}[t]=|a_{z}^{U}[t]|/t$ denotes the vertical direction velocity of the UAV. Moreover, $h[0]$ and $h[T]$ are the UAV flight altitudes at start time and end time, respectively, $m_{U}$ and $g$ denote the mass and gravitational acceleration of the UAV, respectively.
@@ 4.1 Problem Statement
In the considered scenario, the positions of the LEO satellites and vessels are not controllable. Specifically, the vessels sail along specific routes dictated by engines and perform tasks such as sensing and data collection. Moreover, the LEO satellites operate in particular orbits, and the movements are governed by the orbital mechanics. The altitude and inclination of the orbit of the satellite are precisely designed to achieve a specific observation or communication mission.
In this case, this work utilizes the UAV to transmit jamming noise to Eve to resist eavesdropping, thereby enabling secure communications from the LEO satellite to Alice. Due to the spectrum scarcity problem, the UAV and the LEO satellite share the spectrum [42]. The jamming link between the UAV and Eve may interfere with Alice. To address this, we need to regulate the jamming signals transmitted by the UAV, improving its effect on Eve while minimizing the interference on Alice. Therefore, we focus on maximizing the secrecy rate as shown in Eq. (10), which can be controlled by the 3D position and transmit power of the UAV. Note that regulations in the 3D position inevitably result in additional energy consumption of the UAV. Therefore, minimizing the positional regulations of the UAV is essential to enhance overall energy efficiency.
@@ 4.2 Problem Formulation
In the low-altitude friendly-jamming LEO satellite-maritime communication system, we contemplate the following two optimization objectives.
@@
Maximizing the average secrecy rate requires regulating the position of the UAV, which conflicts with minimizing the average energy consumption of the UAV. Moreover, according to Eq. (11), higher UAV speeds lead to increased energy consumption. Conversely, as the UAV slows down, communication time increases, leading to higher hovering energy consumption. Consequently, the two optimization objectives are in conflict, requiring a trade-off.
@@
$\min_{\left\{\mathbb{L},\mathbb{P}\right\}}$ $F=\left\{-f_{1},f_{2}\right\},$ (15a)
@@ 4.3 Problem Analyses
(i) The SSMCMOP is a dynamic problem: In our considered scenario, both the LEO satellites and vessels move along the respective orbits, reflecting a realistic setup. In this case, the UAV needs to transmit jamming signals based on the position of the vessel, leading to a dynamic communication channel. At this point, the SSMCMOP involves two optimization objectives over $\mathcal{T}$ time slots, both of which change in real-time. Therefore, SSMCMOP is a dynamic problem.
@@
$\min_{\left\{\mathbb{L},\mathbb{P}\right\}}$ $F=-f_{1},$ (16a)
@@
Clearly, the simplified SSMCMOP is a classic nonlinear multidimensional 0-1 knapsack problem, which has been proved to be NP-hard [54]. Consequently, the original continuous SSMCMOP is an NP-hard problem.
In summary, the SSMCMOP presents significant challenges, the traditional convex optimization methods and evolutionary computation methods struggle to address the dynamic problem [18]. In this case, the DRL can adaptively learn strategies through continuous environmental interaction and can quickly generate actions to respond to real-time changes [38]. Therefore, we employ a DRL-based algorithm to tackle the SSMCMOP.
@@ 5.1 MDP Formulation
In the considered scenario, our main concern is ensuring the availability of trained DRL models. To this end, we transform the SSMCMOP into an MDP. Specifically, MDP is a mathematical framework used to model decision-making in an uncertain environment, defined by $\left\{\mathcal{S,A,P},\mathcal{\mathbf{R}},\gamma\right\}$ [55], which is introduced in detail as follows.
@@ 5.1.1 State Space
where $\boldsymbol{\eta}_{A}[t]$ and $\boldsymbol{\nu}_{A}[t]$ are the position and velocity of Alice at time slot $t$, respectively. Moreover, $\boldsymbol{\eta}_{E}[t]$ and $\boldsymbol{\nu}_{E}[t]$ are the position and velocity of Eve at time slot $t$, respectively. Note that the current scenario includes a UAV, and our algorithm can be extended to scenarios involving multiple UAVs.
@@ 5.1.3 Immediate Reward
The reward mechanism acts as a feedback signal to direct the agent in a series of actions and directly affects the quality of the final policy. Therefore, designing an effective reward method is critical to improve overall performance. In our work, the MDP employs a reward function in vector form, which is defined by
$\begin{split}\mathcal{\mathbf{R}}&=\left\{\mathcal{R}_{1}(\boldsymbol{s}[t],\boldsymbol{a}[t]),\mathcal{R}_{2}(\boldsymbol{s}[t],\boldsymbol{a}[t])\right\}\\ &=\left\{\begin{array}[]{ll}\left(\mu_{1}R_{SEC}[t],-\mu_{2}E_{U}[t]\right),&k[t]=1,\\ \left(\mu_{1}\varrho_{1}R_{SEC}[t],-\mu_{2}\varrho_{2}E_{U}[t]\right),&k[t]=0,\end{array}\right.\end{split}$ (20)
where $\mathcal{R}_{1}[t]$ and $\mathcal{R}_{2}[t]$ denote the scaled reward values of the optimization objectives $R_{SEC}[t]$ and $E_{U}[t]$, respectively. Moreover, when the UAV moves within the defined range at time slot $t$, $k[t]$ is assigned to 1. Otherwise, $k[t]$ is assigned to 0. Moreover, the coefficients $\mu_{1}$ and $\mu_{2}$ are proportionality factors used to ensure that the rewards for both targets are on the same order of magnitude. Additionally, $\varrho_{1}$ and $\varrho_{2}$ are used to penalize when the UAV leaves the service area.
Following this, we utilize the weighted sum approach to compute the overall expected return [56], which is denoted by

### v2

@@ 3.2 LEO Satellite Orbit Model
LEO satellites orbit at altitudes ranging from about 500 to 2,000 kilometers above the surface of Earth [Deng2021]. Since the orbital altitudes are relatively low, LEO satellites are characterized by fast orbital speeds, usually circling the Earth every 90 to 120 minutes [Pokhrel2024]. Mathematically, the motion of the LEO satellite is usually described by the Keplerian six elements $\left\langle\beta,\omega,\Omega,e,a,\vartheta\right\rangle$ [Deng2021], as shown in the orbit model in Fig. 3, which is introduced in detail as follows.
@@
We consider that the LEO satellite orbit is circular [Deng2021]. In this case, $e$ and $\vartheta$ are set to 0, and $a$ is equal to the orbital radius $l_{S}$. Accordingly, the elements of the LEO satellite in $m$ orbit at $t$ time can be set as $\left\langle\beta_{m},\omega_{m}[t],\Omega_{m},{l}_{Sm}\right\rangle$, wherein $\omega_{m}[t]=\omega_{m}[t]+2\pi\left(({t}/{\mathcal{T}_{Sm}})\mod 1\right)$ and $\mathcal{T}_{Sm}$ denotes the orbital period of the LEO satellite. Moreover, $l_{Sm}=H_{Sm}+R_{E}$, where $H_{Sm}$ and $R_{E}$ are the orbital altitude of the LEO satellite and the radius of Earth, respectively. In this case, the orientation of the LEO satellite in the 3D Cartesian coordinates $(x_{Sm}[t],y_{Sm}[t],z_{Sm}[t])$ at time slot $t$ is expressed by
@@ 3.3 Vessel Movement Model
As shown in Fig. 3, two three-dimensional right-handed Cartesian coordinate systems of a vessel are generally used to represent movement of the vessel [Skulstad2020]. Specifically, one is a general coordinate system $\left\{g\right\}$ in which the origin is on the surface of the sea, and the defined $x$, $y$, and $z$ axes are pointing the north, east, and down, respectively. The other is a fixed coordinate system $\left\{f\right\}$ in which the origin is the gravity of the vessel, with $x^{f}$, $y^{f}$, and $z^{f}$ pointing to the bow, starboard, and down, respectively. The rotations around the $x^{f}$, $y^{f}$, and $z^{f}$ axes are defined as roll ($\phi$), pitch ($\theta$), and yaw ($\psi$), respectively. Moreover, the Euler angle vector is denoted as $\boldsymbol{\Theta}=\left[\phi,\theta,\psi\right]^{T}$. Mathematically, the six degree-of-freedom (DOF) vessel model $\left\langle x,y,z,\phi,\theta,\psi\right\rangle$ is used to represent movement of a vessel, which is as follows:
@@ 3.4.1 S2V Link from LEO Satellite
During the S2V link, we consider that the vessels can detect the signals from the LEO satellite by using the approach in [Guo2022], thus obtaining the quantized form of the actual channel state information (CSI). In this case, since the altitude of the LEO satellite is sufficient for line-of-sight (LoS) transmission [Shang2024], we employ a typical composite channel to represent the communications from the LEO satellite to vessels. The channel of the S2V link at time slot $t$ can be expressed by [Li2020]
@@
where $C_{S}$ and $W_{S}$ are the path loss parameter and exponent, respectively. Moreover, $d_{S,V}[t]$} is the distance between the LEO satellite and vessel (Alice or Eve) at time slot $t$, which is computed by the 3D positions of the LEO satellite and vessel according to Eqs. (1) and (2), respectively. In addition, $\delta_{S,V}[t]$ is the zero-mean Gaussian random variable with standard deviation $\sigma_{X_{S}}$ [Wang2018][Wu2018a]. Note that $d_{S,A}[t]$ and $d_{S,E}[t]$ are the distance from the LEO satellite to Alice and Eve at time slot $t$, respectively, which are used to calculate corresponding path loss ($PL_{S,A}[t]$ and $PL_{S,E}[t]$), and get the matching channels ($h_{S,A}[t]$ and $h_{S,E}[t]$).
@@ 3.4.2 U2V Link from UAV
where $d_{U,V}[t]$} denotes the distance from the UAV to vessel (Alice or Eve) at time slot $t$, which is computed based on the movements of the UAV and vessel. Moreover, $C_{U}$ denotes the path loss parameter, $d_{c}$ denotes the reference distance, $W_{U}$ and $\delta_{U,V}[t]$ are the path loss parameter and zero-mean Gaussian random variable with $\sigma_{X_{U}}$, respectively [Matolak2016]. Note that $PL_{U,A}[t]$ and $PL_{U,E}[t]$ denote the path loss from the UAV to Alice and Eve at time slot $t$, respectively, which are obtained by the distance from the UAV to Alice and Eve ($d_{U,A}[t]$ and $d_{U,E}[t]$), respectively.
@@
where $P_{{S}}$ is the transmit power of the LEO satellite, $G_{S}$ and $G_{U}$ denote the antenna gains of the satellite and UAV, respectively. Moreover, $G_{S,S}$ is the antenna gain of the vessel served by the LEO satellite, $P_{{U}}[t]$ denotes the transmit power of the UAV at time slot $t$, and $\sigma^{2}$ denotes the maritime additive white Gaussian noise power.
@@
where $\left[\chi\right]^{+}$ indicates the value at which the larger of 0 and $\chi$.
@@ 3.5 UAV Energy Consumption Model
Furthermore, we introduce the moving UAV energy consumption. Commonly, the total energy consumption of a UAV is composed of communication energy and propulsion energy. However, the communication energy is usually ignored during the calculation because its value is extremely small compared to propulsion energy [Zeng2018]. Thus, the propulsion power consumption of the UAV in 2D horizontal state is calculated by [Zeng2018]
@@
where $v_{h}[t]=\sqrt{(a_{x}^{U}[t])^{2}+(a_{y}^{U}[t])^{2}}/\Delta t$ is the horizontal direction velocity of the UAV, $P_{I}$ and $P_{B}$ are the induced power and blade profile power, respectively, $v_{tip}$ and $v_{i}$ are the rotor blade tip velocity and average rotor induced velocity, respectively. Moreover, $r_{a}$, $s_{r}$, $a_{r}$, and $\rho$ denote the airframe drag ratio, rotor robustness, rotor disk area, and air density, respectively.
Note that we exclude the energy consumed by the UAV accelerating or decelerating, since this process occupies only a minor fraction of the total UAV flight time [Li2023a]. Therefore, we use a heuristic closed approximation to express the 3D energy consumption of the UAV, where the considerations include the propulsive energy, kinetic energy, and gravitational energy during the ascent and descent of the UAV over time. The 3D trajectory energy consumption of the UAV is denoted by
@@
where $v_{f}[t]=\sqrt{\left(v_{h}[t]\right)^{2}+\left(v_{v}[t]\right)^{2}}$ is the UAV forward direction velocity at time slot $t$, of which $v_{v}[t]=|a_{z}^{U}[t]|/\Delta t$ denotes the vertical direction velocity of the UAV. Moreover, $h[0]$ and $h[T]$ are the UAV flight altitudes at start time and end time, respectively, $m_{U}$ and $g$ denote the mass and gravitational acceleration of the UAV, respectively.
@@ 4.1 Problem Statement
In the considered scenario, the trajectories of the LEO satellites and vessels are not controllable. Specifically, the vessels sail along specific routes dictated by engines and perform tasks such as sensing and data collection. Moreover, the LEO satellites operate in particular orbits, and their movements are governed by orbital mechanics. The altitude and inclination of the orbit of the satellite are precisely designed to achieve a specific observation or communication mission.
In this case, this work utilizes the UAV to transmit jamming noise to Eve to resist eavesdropping, thereby enabling secure communications from the LEO satellite to Alice. Due to the spectrum scarcity problem, the UAV and the LEO satellite share the spectrum [Li2020]. The jamming link between the UAV and Eve may interfere with Alice. To address this, we need to regulate the jamming signals transmitted by the UAV, improving its effect on Eve while minimizing the interference on Alice. Therefore, we focus on maximizing the secrecy rate as shown in Eq. (10), which can be controlled by the 3D position and transmit power of the UAV. Note that regulations in the 3D position inevitably result in additional energy consumption of the UAV. Therefore, minimizing the positional regulations of the UAV is essential to enhance overall energy efficiency.
@@ 4.2 Problem Formulation
In the LEO satellite-maritime communication system assisted by a low-altitude UAV, we contemplate the following two optimization objectives.
@@
Maximizing the average secrecy rate requires regulating the position of the UAV, which conflicts with minimizing the average energy consumption of the UAV. Moreover, according to Eq. (11), higher UAV speeds lead to increased energy consumption. Conversely, as the UAV slows down, communication time increases, leading to higher hovering energy consumption. Consequently, the two optimization objectives are in conflict, requiring a suitable modeling method to balance this conflicting relationship. In this case, the multi-objective optimization problem modeling provides a mathematical framework that simultaneously optimizes multiple conflicting objectives, which is well-suited for capturing trade-offs of conflicting metrics and can be used to formulate our optimization problem [Karami2022].
@@
$\min_{\left\{\mathbb{L},\mathbb{P}\right\}}\$$F=\left\{-f_{1},f_{2}\right\},$ (15a)
@@ 4.3 Problem Analyses
(i) The SSMCMOP is a dynamic and large-scale problem: In our considered scenario, both the LEO satellites and vessels move along the respective orbits, reflecting a realistic setup. In this case, the UAV needs to transmit jamming signals based on the position of the vessel, leading to a dynamic communication channel. At this point, the SSMCMOP involves two optimization objectives over $\mathcal{T}$ time slots, both of which change in real-time. In addition, the UAV has multiple variables (3D position and transmit power) to be optimized. Therefore, SSMCMOP is a dynamic and large-scale problem.
@@
$\min_{\left\{\mathbb{L},\mathbb{P}\right\}}\$$F=-f_{1},$ (16a)
@@
Clearly, the simplified SSMCMOP is a classic nonlinear multidimensional 0-1 knapsack problem, which has been proved to be NP-hard [Goos2020]. Consequently, the original continuous SSMCMOP is an NP-hard problem.
In summary, since the SSMCMOP presents significant challenges, traditional convex optimization methods and evolutionary computation methods struggle to address the dynamic problem [Wang2023]. In this case, the DRL algorithms can adaptively learn strategies through continuous environmental interaction in the online training phase, followed by offline execution where the trained model can quickly generate actions to respond to real-time changes [Li2024]. Therefore, we employ a DRL-based algorithm to tackle the SSMCMOP.
@@ 5.1 MDP Formulation
In the considered scenario, our main concern is ensuring the availability of trained DRL models. To this end, we transform the SSMCMOP into an MDP. Specifically, MDP is a mathematical framework used to model decision-making in an uncertain environment, defined by $\left\{\mathcal{S,A,P},\mathcal{\mathbf{R}},\gamma\right\}$ [Xu2020], which is introduced in detail as follows.
@@ 5.1.1 State Space
where $\boldsymbol{\eta}_{A}[t]$ and $\boldsymbol{\nu}_{A}[t]$ are the position and velocity of Alice at time slot $t$, respectively. Moreover, $\boldsymbol{\eta}_{E}[t]$ and $\boldsymbol{\nu}_{E}[t]$ are the position and velocity of Eve at time slot $t$, respectively. Note that the current scenario includes a UAV with a single antenna, which limits the UAV to jam one illegitimate vessel at a time. Furthermore, our algorithm can be extended to scenarios with multiple UAVs, and the details are provided in Section 7.
@@ 5.1.3 Immediate Reward
The reward mechanism acts as a feedback signal to direct the agent in a series of actions and directly affects the quality of the final policy. Therefore, designing an effective reward method is critical to improve overall performance. Note that the constraints $C1$-$C5$ of the SSMCMOP are satisfied by setting the parameters of the UAV within the predetermined ranges. Moreover, the constraints $C6$ and $C7$ are satisfied by incorporating them into the reward function as penalty items, and we further incorporate constraints $C1$, $C2$, and $C3$ into the reward function to ensure strict adherence to flight boundaries. In our work, the MDP employs a reward function in vector form, which is defined by
[…]
where $\mathcal{R}_{1}[t]$ and $\mathcal{R}_{2}[t]$ denote the scaled reward values of the optimization objectives $R_{SEC}[t]$ and $E_{U}[t]$, respectively. Moreover, when the UAV moves within the defined range at time slot $t$, $k[t]$ is assigned to 1. Otherwise, $k[t]$ is assigned to 0. Moreover, the coefficients $\mu_{1}$ and $\mu_{2}$ are proportionality factors used to ensure that the rewards for both targets are on the same order of magnitude. Additionally, $\varrho_{1}$ and $\varrho_{2}$ are used to penalize when the UAV leaves the service area. Additionally, $W_{c}[t]=W_{1}[t]W_{2}[t]$ represents the comprehensive penalty term of the constraints $C6$ and $C7$ of the SSMCMOP. Correspondingly, $W_{1}$ and $W_{2}$ represent the penalty items when the strength of the interference signals received by Alice and Eve affects the maritime communications, respectively.
Following this, we utilize the linear weighting approach to compute the overall expected return [Marler2010]. This approach offers low computational complexity and adjustable priority, making it well-suited for the SSMCMOP in resource-constrained dynamic maritime environments. The overall expected return is denoted by